# A classification result for eternal mean convex flows of finite total curvature type

### 18 Mar 2024 v1

@@ 4. Concluding remarks
First, we further discuss the case left from above that $M_{t}\simeq{\mathbb{R}}^{2}$. Under the finite total curvature type condition this set is nonempty, considering the bowl soliton, but there are some hints to suggest there are none under a finite total curvature assumption of the form “there is some constant $C>0$ so that $\int_{M_{t}}|A|^{2}<C$ for all $t<0$.” With this assumption (or one that implies this) to proceed it seems natural to consider the blowdown of $M_{t}$. By this we recall one parabolically rescales $M_{t}$ about the origin by a sequence of scales tending $\lambda_{i}\to 0$. Denoting the resulting flows by $M^{\lambda_{i}}_{t}$, by the finite entropy assumption as in the section above one may employ Brakke’s compactness theorem (considering the flows as Brakke flows) to extract a limiting Brakke flow which, by Huisken’s monotonicity formula, must be a self shrinker. As discussed in [16], its support will be smooth although the convergence may be with multiplicity greater than one and there are examples where this is the case: for instance, considering the catenoid as an eternal flow whose blowdown is the plane with multiplicity 2.
@@
Since $M_{t}$ is mean convex the blowdown limit as well must be mean convex, so by the finite entropy assumption and [5] must either be a flat plane or round cylinder (or round sphere, which is ruled out by the noncompactness of $M_{t}$). In the later case, using that the total curvature is scale invariant we find a contradiction since the round cylinder has infinite total curvature (in terms of integral of $|A|^{2}$, not $|K|$ of course). The issue, at least from this line of attack, is if the blowdown is a plane. Indeed, its blowdown limit cannot be a flat plane of multiplicity one – if this were the case then by Huisken’s monotonicty the flow $M_{t}$ itself must be a flat plane, contradicting that we assume $H>0$ strictly along the flow. It seems that ruling out the case the blowdown limit is a plane of multiplicity 2 or higher though is more delicate, but to the author’s knowledge the known examples of such flows diffeomorphic to the plane, such as the grim reaper cylinder and the $\Delta$–wings (see [13, 2]), have infinite total curvature. There are also more exotic convex eternal flows which aren’t translators, constructed in [1].
@@
Note that in the above no claims of uniqueness are made for the eternal solution up to the catenoid it flows out of – it seems quite plausible that such a solution is unique though. For instance, the authors showed in [4] that there was only one ancient rescaled mean curvature flow which layed to one side (apriori not even shrinker mean convex) of an asymptotically conical self shrinker, using the Merle–Zaag ODE lemma. There are a number of issues one needs to address though, for instance how well an eternal solution $M_{t}$ can be shown to converge back in time to its catenoid $C$ – consider lemma 7.18 in [4]. We also mention that in [22] some uniqueness results for suitably symmetric eternal mean convex flows were shown – see theorem 1.5 therein. In a nutshell, the assumptions quickly imply that such a flow must asymptote to a catenoid as $t\to-\infty$, and using uniqueness results for the curve shortening flow we can show an “asymptotic uniqueness” statement.
@@
As discussed already in the introduction, if we relax our hypotheses of mean convexity we then must include all minimal surfaces of fiinte topology, the space of which is comparitively quite large with varied topology. In this space there may also be topologically interesting translators of finite total curvature. By [18] such translators must have asymptotically planar ends and with this in mind one could imagine, for instance, the existence of a “sideways translating annulus” of finite total curvature, although these have recently been conjectured to not exist [12]. If we consider relaxing our assumption of finite total curvature type there are clearly weaker assumptions which will allow us to take a backwards limit and get something smooth, although possibly not of finite total curvature depending on the manner the assumptions were loosened which played a role in the argument. There is also the question of generalizing this result to higher dimensions, which is reasonable to ask because the reapernoid of [22] was constructed for all $n\geq 2$, although we point out that in a number of steps (for instance in taking the backwards limit) that we were in the surface case was strongly used.

### 14 Mar 2026 v2

@@ 4. Concluding remarks
First, we further discuss the case left from above that $M_{t}\simeq{\mathbb{R}}^{2}$. Under the finite total curvature type condition this set is nonempty, considering the bowl soliton, but there are some hints to suggest there are none under a finite total curvature assumption of the form “there is some constant $C>0$ so that $\int_{M_{t}}|A|^{2}<C$ for all t<0t<0.” With this assumption (or one that implies this) to proceed it seems natural to consider the blowdown of $M_{t}$. By this we recall one parabolically rescales $M_{t}$ about the origin by a sequence of scales tending $\lambda_{i}\to 0$. Denoting the resulting flows by $M^{\lambda_{i}}_{t}$, by the finite entropy assumption as in the section above one may employ Brakke’s compactness theorem (considering the flows as Brakke flows) to extract a limiting Brakke flow which, by Huisken’s monotonicity formula, must be a self shrinker. As discussed in [16], its support will be smooth although the convergence may be with multiplicity greater than one and there are examples where this is the case: for instance, considering the catenoid as an eternal flow whose blowdown is the plane with multiplicity 2.
@@
Since $M_{t}$ is mean convex the blowdown limit as well must be mean convex, so by the finite entropy assumption and [5] must either be a flat plane or round cylinder (or round sphere, which is ruled out by the noncompactness of $M_{t}$). In the later case, using that the total curvature is scale invariant we find a contradiction since the round cylinder has infinite total curvature (in terms of integral of $|A|^{2}$, not $|K|$ of course). The issue, at least from this line of attack, is if the blowdown is a plane. Indeed, its blowdown limit cannot be a flat plane of multiplicity one – if this were the case then by Huisken’s monotonicity the flow $M_{t}$ itself must be a flat plane, contradicting that we assume $H>0$ strictly along the flow. It seems that ruling out the case the blowdown limit is a plane of multiplicity 2 or higher though is more delicate, but to the author’s knowledge the known examples of such flows diffeomorphic to the plane, such as the grim reaper cylinder and the $\Delta$–wings (see [13, 2]), have infinite total curvature. There are also more exotic convex eternal flows which aren’t translators, constructed in [1].
@@
Note that in the above no claims of uniqueness are made for the eternal solution up to the catenoid it flows out of – it seems quite plausible that such a solution is unique though. For instance, it is shown in [4] that there was only one ancient rescaled mean curvature flow which laid to one side (a priori not even shrinker mean convex) of an asymptotically conical self shrinker, using the Merle–Zaag ODE lemma. There are a number of issues one needs to address though, for instance how well an eternal solution $M_{t}$ can be shown to converge back in time to its catenoid $C$ – consider lemma 7.18 in [4]. We also mention that in [22] some uniqueness results for suitably symmetric eternal mean convex flows were shown – see theorem 1.5 therein. In a nutshell, the assumptions quickly imply that such a flow must asymptote to a catenoid as $t\to$-$\infty$, and using uniqueness results for the curve shortening flow we can show an “asymptotic uniqueness” statement.
@@
As discussed already in the introduction, if we relax our hypotheses of mean convexity we then must include all minimal surfaces of finite topology, the space of which is comparatively quite large with varied topology. In this space there may also be topologically interesting translators of finite total curvature. By [18] such translators must have asymptotically planar ends and with this in mind one could imagine, for instance, the existence of a “sideways translating annulus” of finite total curvature, although these have recently been conjectured to not exist [12]. If we consider relaxing our assumption of finite total curvature type there are clearly weaker assumptions which will allow us to take a backwards limit and get something smooth, although possibly not of finite total curvature depending on the manner the assumptions were loosened which played a role in the argument. There is also the question of generalizing this result to higher dimensions, which is reasonable to ask because the reapernoid of [22] was constructed for all $n\geq 2$, although we point out that in a number of steps (for instance in taking the backwards limit) that we were in the surface case was strongly used.